OpenStudy (anonymous):

Calculate that : \[\int\limits_{0}^{3}((2x+3)/(x+1))dx + \ln3\]

OpenStudy (lalaly):

\[\int\limits{\frac{2x+3}{x+1}dx}= \int\limits{\frac{1}{x+1}dx} + \int\limits{3dx} +\ln3\]

OpenStudy (anonymous):

did not get :(

OpenStudy (lalaly):

long division